# Measuring the splashback feature: Dependence on halo properties and history

### v1

@@ 2.5 The splashback feature
Figure 1: The upper panel is the sample density profile of the dark matter halo, and the lower panel is the logarithmic slope of the density profile. The data points of the density profile come from the MTNG hydrodynamic simulations. This example is for the mass range $10^{14}\leq M/{\rm M}_{\odot}<10^{14.5}M_{\odot}$. The vertical black dashed line indicates the location of steepest slope $R_{\rm st}$ of the stacked haloes. The solid blue line is the fitted density profile and logarithmic slope using Equation 7 and Equation 10, respectively. In the lower panel, the horizontal green and the vertical red lines indicate the width and the depth of the splashback features defined in Section 2.5.
The slope of the halo density profiles has a valley shape. Apart from the position of the steepest slope, the width and depth of the logarithmic slope valley also track the kinematics at the halo boundaries. Therefore, in order to provide quantitative measures, we define the splashback depth $\mathcal{D}$ as the difference between the logarithmic slope at the splashback radius and the radius where the gradient profile turns over outside the halo. This is the point of maximum curvature in the gradient to the right of the minimum. Mathematically, the splashback depth is defined as
@@
Then the width $\mathcal{W}$ of the logarithmic slope valley is the width of the feature halfway between the height and the minimum of the feature. A visualisation of the splashback depth and width is shown in Figure 1. These two features relate to the splashback radius and the overall halo density profiles. Although the depth and the width can possibly correlate with each other, we treat them as separate variables and examine how they vary with halo properties over time. We also examine the ratio of the two in some cases to characterize the overall sharpness of the feature.
@@ 3.1 Splashback features as a function of halo mass and redshift
Figure 2: Depth (Left) and width (Right) of splashback features as a function of halo mass for the dark matter mass at $z=0$. We stack the density profiles for haloes with $\log\left(M_{200\text{m}}/M_{\odot}\right)$ in 13–13.5, 13.5–14, 14–14.5, 14.5–15, and 15-15.5 and compute the splashback features of the median profile. The depth increases with mass while the width decreases. This indicates a more pronounced splashback feature for larger haloes. Notably, this is not just an enlargement of the feature, in which both the width and depth would increase, but the decreasing width indicates a sharper decrease in the density profile.
Figure 3: Top left: Depth of splashback features as a function of halo mass for the dark matter mass at $z<2$. The depth linearly increases with the log of the mass. Top right: Depth of splashback features as a function of redshift for the dark matter halo mass with $13\leq\log\left(M_{200\text{m}/M_{\odot}}\right)\leq 15.5$. There is also a slight increasing trend with redshift, although this is much less pronounced than the trend with depth. Bottom left: width of splashback features as a function of halo mass for the dark matter mass at $z<2$. The width decreases with mass, indicating that the splashback feature is narrower for larger mass haloes. Bottom right: Width of splashback features as a function of redshift for the dark matter halo mass with $13\leq\log\left(M_{200\text{m}/M_{\odot}}\right)\leq 15.5$. There is a decreasing trend with redshift, indicating that the width is more sensitive to cosmic evolution than the depth. The dashed lines are referred to the power law fitting in Equations 3.1 and 3.1.
We start with examining the splashback depth and width as a function of halo mass at redshift $z=0$. As indicated in Section 2.3, we focus on the hydrodynamic simulation in MTNG and split the haloes into 5 mass bins between $10^{13}$ and $10^{15.5}M_{\odot}$. We calculate the depth and width of the splashback features of the stacked haloes in each bin by applying the fitting methods introduced in Section 2.4. Figure 2 shows the change of splashback features as a function of mass at $z=0$. The depth of the splashback feature increases steadily from 2.2 to 3.0 with mass, and the width decreases from 2.0 to 1.0. The splashback features of haloes become more significant when the halo mass is more massive. The same computations are also performed on the dark matter-only simulation in MTNG and the splashback features calculated are shown in Appendix B. There is no significant difference between the hydrodynamic and dark matter-only simulations so we will focus on the hydrodynamic simulation only in following sections.
@@
To generalise this expression, we examine how the splashback features change as a function of mass at different redshifts. We trace the redshift back to $z=2$ and calculate the splashback features at each redshift. The result is shown in the left panels of Figure 3. At higher redshifts, the depth of the splashback feature consistently increases with halo mass, while its width consistently decreases. This trend therefore persists across redshifts. It indicates that more massive haloes possess clearer boundaries and a clearer halo boundary would lead to deeper and narrower splashback features by definition of $\mathcal{D}$ and $\mathcal{W}$.
Based on the power law fitting in Equations 3.1 and 3.1, we add an extra scale factor $a=1/{(z+1)}$ dependence to the width and get the following fitting equations of depth and width of splashback features as a function of both mass and redshift (depth: $\chi_{\nu}^{2}=1.4$; width: $\chi_{\nu}^{2}=2.2$).
@@
Based on the existing fitting equations, we plot the splashback features predicted by halo mass and redshift, which are the dashed lines in the right panels of Figure 3. For haloes with the same mass, the splashback feature is narrower in the past. In addition, the depth does not change significantly with time for a given mass, indicating that the splashback feature is more significant in the past. The narrower and deeper splashback features mean a more rapid change of density of halo in the splashback regions. A possible explanation is that as a halo grows in size with time, by keeping the mass constant, the halo average density is smaller since particles are more spread out. The spreading of particles increases the splashback width. The halo accretion rate also decreases with time (McBride, Fakhouri & Ma, 2009). Slow accretion results in a more gradual transition between virialised and infalling material, thus the splashback features are smeared out. This trend is further verified quantitatively in the next section through a comparison of the splashback features with different accretion rates.
Qualitatively, we observe that the depth has a smaller dependence on the redshift compared with the width, since the former is approximately constant for different redshift and the latter is spread out with redshift. This indicates that the splashback width is more affected by the cosmic evolution than the depth is, but overall haloes with the same mass have more distinct splashback features at higher redshifts.
@@ 3.2 Splashback features as a function of halo accretion rate and merger history
Figure 5: Relationships between width ($\mathcal{W}$) and depth ($\mathcal{D}$) of the splashback feature as a function of redshift ($z$) and accretion rate ($\Gamma$). When haloes are stacked by accretion rate, the width increases with time. However, the depth is uncorrelated with redshift and both width and depth appear only weakly dependent on accretion rate. In all panels, Points show the data from fits in the simulation, with colour indicating either redshift (left) or accretion rate (right). Haloes are stacked into accretion rate bins between $\Gamma=0$ and $\Gamma=6$ and the median profile is fitted. Top left: Depth of the splashback features as a function of accretion rate for various redshifts. Top right: Depth of the splashback feature as a function of redshift for various accretion rate. There is no a clear trend of the depth and width as a function of accretion rate. Bottom left: Width of the splashback features as a function of accretion rate for various redshifts. Bottom right: Width of the splashback feature as a function of redshift for various accretion rates. The black dashed line is fitted with the redshift only while the black solid line is fitted using the scale factor. There is not much of a trend for the depth and width in terms of halo accretion rate, so we fit a function only in terms of redshift. There is less dependence on redshift when separated by accretion rate compared to the mass bins, as shown in Figure 3, although there is still some dependence with the depth decreasing significantly between redshift $0<z<1$ and with the width peaking at redshift $z\sim 1$.
@@
[…]
Several studies (e.g. Diemer & Kravtsov, 2014; Adhikari, Dalal & Chamberlain, 2014) show that the splashback feature has a strong dependence on the halo mass accretion rate. Diemer & Kravtsov (2014) notes that high accretion rate haloes with $\Gamma\gtrsim 3$ have a logarithmic slope that reach approximately 4-4- 4 while lower accretion rate haloes with $\Gamma\lesssim 1$ have logarithmic slopes that reach only 3-3- 3. This observation correlates with the splashback depth defined in our study.
@@
We start with haloes at $z=0$. In Figure 4, the splashback depth gradually increases from 2.4 to 3.0 with the accretion rate from $\Gamma=0$ to $\Gamma=5$. The result matches the observations and conclusions in Diemer & Kravtsov (2014). Although the definition of the splashback depth is not the same as the logarithmic slope minimum in Diemer & Kravtsov (2014), they demonstrate a similar trend. We observe that, at $\Gamma\approx 1$, the splashback depth is approximately $\mathcal{D}=2.4\pm 0.1$; at $\Gamma\approx 3$, the splashback depth is around $\mathcal{D}=2.5\pm 0.1$. This value is consistent with that that found in Diemer & Kravtsov (2014). Assuming our simulations match the result in Diemer & Kravtsov (2014), this consistency predicts that, at higher halo accretion rate, around $\Gamma=5$, the splashback depth is 3.0±0.63.0$\pm$ 0.63.0 ± 0.6. We attempt to apply the numerical fitting to the splashback depth $\mathcal{D}$ in terms of halo accretion rate. We implement a power law and a linear fitting as below
@@
and they respectively possess reduced $\chi_{\nu}^{2}=0.10$ and $\chi_{\nu}^{2}=0.15$. It is unclear which fitting method better describes the data. One possible explanation for the power-law dependence of splashback depth on peak height is its indirect connection to halo accretion rate. As shown in Diemer & Kravtsov (2014), the accretion rate increases approximately linearly with peak height. In the following section, we show that the splashback depth scales as a power law with peak height, suggesting that the depth may inherit this dependence through the accretion history of the halo.
@@
Therefore, faster accreting haloes have more pronounced halo boundaries and more distinct splashback features. The decrease of the splashback width is also correlated with the results found in Figure 2. Since the halo accretion rate $\Gamma$ increases with the halo mass $M_{200\text{m}}$ and peak height, the power law decrease of the splashback width with accretion rate is consistent with that found in Section 2.2.1 and also following Section 3.3.2. As haloes accrete matter more rapidly, newly infalling particles pile up more tightly near the first apocentre, leading to a steeper density gradient and a narrower transition zone between the virialised and infalling regions. This results in a narrower splashback width and a more pronounced splashback feature.
In order to check if the splashback feature as a function of accretion rate also persists at higher redshifts, we go back to $z=2$ and plot the depth and width in terms of accretion rate and redshift in Figure 5. At higher redshift, the splashback features tend to be more independent of the halo accretion rate and the previous power law fitting fails. The depth varies between $\mathcal{D}=2.4\sim 2.8$, and the width remains constant between $1\leq\Gamma\leq 6$. At $z\sim 2$ and $\Gamma\gtrsim 4$, the splashback depth decreases with accretion rate. By examining the halo profiles at $z=2$, this can be explained because faster accreting haloes have more turbulent environments (i.e. more density fluctuations at the edges) (Klessen & Hennebelle, 2010), which causes the outskirts past $R_{\rm st}$ to flatten to a lower value at large radius. Meanwhile, the width at higher redshifts remains insensitive to accretion rate, indicating that the halo boundary remains largely unaffected by accretion rate. Overall, the accretion rate at high redshift affects the halo environment more than the halo internal density structure.
To account for this difference, we examine the halo mass-accretion rate distributions at $z=0$ and $z=2$, which are shown in Figure 6. It shows at, at present, among different accretion bins, the halo mass distributions are different. At $z=0$, the median halo mass in each accretion bin is weakly correlated with accretion rate. In contrast, at $z=2$, the median halo mass remains approximately constant across accretion bins. This may be partially due to the lack of large haloes at $z=2$ making the trend less significant at higher redshift. Therefore, when bootstrapping haloes in each accretion rate bin, there is a stronger dependence on the halo mass at $z=0$ and so higher similarity with splashback features binned by mass in Figure 3. On the other hand, as the mass distributions in each accretion bin are similar at $z=2$, this removes the mass dependence during bootstrapping.
@@
with $\chi_{\nu}^{2}=2.4$. The constant $C$ is the splashback width at $z=0$. More specifically, it is a halo mass accretion rate-dependent variable $C(\Gamma)$ at the present, which is effectively Equation 17 and 19. The width can therefore be fitted as
@@
Unrelaxed haloes that are still adjusting post-merger may exhibit disrupted or less refined splashback features due to chaotic particle orbits compared to relaxed haloes with more stable structures. More et al. (2016) noted that the variations in $\frac{R_{\rm sp}}{R_{\rm 200m}}$ due to variations in relaxation state is approximately 0.2. Therefore, we define the halo merging time in Section 2.2.3 to track this effect. The halo merges when the ratio between the halo’s second most massive subhalo and the most massive subhalo is greater than 0.1. We focus on the haloes at redshift $z=0$. We select haloes with a major merging time less than $z=9$ and and stack the haloes into bins with an interval of $\Delta z=0.5$. In the upper panel of Figure 7, the splashback depth increases before $z=3$, and tends to stability afterwards. This verifies that relaxed haloes which merged at higher redshift tend to have more stable splashback features compared with unrelaxed haloes which merge at $z<3$. The splashback width is relatively constant at $z<6$ due to large error bars. At $z>6$, the splashback width reduces slightly for $\Delta W\approx 0.1$.
@@
However, the overall density profile is the same for haloes with both recent major mergers and haloes with major mergers that occurred longer ago. Therefore, the splashback feature depth-to-mass ratio is independent of the halo merging time. One thing to note is that these are stacked profiles, but the mergers occur in individual haloes and at only one point on the halo’s edge. Thus, the effects of mergers could be washed out by stacking, which by design removes spherical asymmetries in the halo profiles. It may be possible to align the stacked haloes by merger direction or stack individual objects by solid angle to better capture these effects, but we leave this for future work.
The splashback features demonstrate less dependence on accretion rate and merging time which are short term factors. This shows that instead of being similar to the splashback radius which has stronger correlation with halo accretion rate, the splashback depth and width depend more on the long term features of haloes. This is possibly due to the depth and width being measures of halo density profile rather than the tracker of particles’ first turnaround at the halo boundary.

### v2

@@ 2.5 The splashback feature
Figure 1: The upper panel is the sample density profile of the dark matter halo, and the lower panel is the logarithmic slope of the density profile. The data points of the density profile come from the MTNG hydrodynamic simulations. This example is for the mass range $10^{14}\leq M/{\rm M}_{\odot}<10^{14.5}M_{\odot}$. The vertical black dashed line indicates the location of the steepest slope $R_{\rm st}$ of the stacked haloes. The solid blue line is the fitted density profile and logarithmic slope using Equation 7 and Equation 10, respectively. In the lower panel, the horizontal green and the vertical red lines indicate the width and the depth of the splashback features defined in Section 2.5.
The slope of the halo density profiles has a valley shape. Apart from the position of the steepest slope, the width and depth of the logarithmic slope valley also track the kinematics at the halo boundaries. Therefore, in order to provide quantitative measures, we define the splashback depth $\mathcal{D}$ as the difference between the logarithmic slope at the splashback radius and the radius where the gradient profile turns over outside the halo. This is the point of maximum curvature in the gradient to the right of the minimum. We compute the curvature of the slope profile in a separate procedure. We first smooth the profile with a one-dimensional Gaussian filter and then apply standard finite-difference derivatives (using np.gradient) to obtain the first, second, and third derivatives. The point of maximum curvature is identified from the extremum of the third derivative. Mathematically, the splashback depth is defined as
@@
Then, the width $\mathcal{W}$ of the logarithmic slope valley is the width of the feature halfway between the height and the minimum of the feature. A visualisation of the splashback depth and width is shown in Figure 1. These two features relate to the splashback radius and the overall halo density profiles. Although the depth and the width can possibly correlate with each other, we treat them as separate variables and examine how they vary with halo properties over time. We also examine the ratio of the two in some cases to characterise the overall sharpness of the feature.
@@ 3.1 Splashback features as a function of halo mass and redshift
Figure 2: Depth (Left) and width (Right) of splashback features as a function of halo mass for the dark matter mass at $z=0$. We stack the density profiles for haloes with $\log\left(M_{200\text{m}}/M_{\odot}\right)$ in 13–13.5, 13.5–14, 14–14.5, 14.5–15, and 15–15.5 and compute the splashback features of the median profile. The depth increases with mass while the width decreases. This indicates a more pronounced splashback feature for larger haloes. Notably, this is not just an enlargement of the feature, in which both the width and depth would increase, but the decreasing width indicates a sharper decrease in the density profile.
Figure 3: Top left: Depth of splashback features as a function of halo mass for the dark matter mass at $z<2$. The depth linearly increases with the log of the mass. Top right: Depth of splashback features as a function of redshift for the dark matter halo mass with $13\leq\log\left(M_{200\text{m}/M_{\odot}}\right)\leq 15.5$. There is also a slight increasing trend with redshift, although this is much less pronounced than the trend with depth. Bottom left: width of splashback features as a function of halo mass for the dark matter mass at $z<2$. The width decreases with mass, indicating that the splashback feature is narrower for larger mass haloes. Bottom right: Width of splashback features as a function of redshift for the dark matter halo mass with $13\leq\log\left(M_{200\text{m}/M_{\odot}}\right)\leq 15.5$. There is a decreasing trend with redshift, indicating that the width is more sensitive to cosmic evolution than the depth. The dashed lines refer to the power law fitting in Equations 3.1 and 3.1.
We start with examining the splashback depth and width as a function of halo mass at redshift $z=0$. As indicated in Section 2.3, we focus on the hydrodynamic simulation in MTNG and split the haloes into 5 mass bins between $10^{13}$ and $10^{15.5}M_{\odot}$. We calculate the depth and width of the splashback features of the stacked haloes in each bin by applying the fitting methods introduced in Section 2.4. Figure 2 shows the change of splashback features as a function of mass at $z=0$. The depth of the splashback feature increases steadily from 2.2 to 3.0 with mass, and the width decreases from 2.0 to 1.0. The splashback features of haloes become more significant when the halo mass is more massive. The same computations are also performed on the dark matter-only simulation in MTNG, and the splashback features calculated are shown in Appendix B. There is no significant difference between the hydrodynamic and dark matter-only simulations so we will focus on the hydrodynamic simulation only in following sections.
@@
To generalise this expression, we examine how the splashback features change as a function of mass at different redshifts. We trace the redshift back to $z=2$ and calculate the splashback features at each redshift. The result is shown in the left panels of Figure 3. At higher redshifts, the depth of the splashback feature consistently increases with halo mass, while its width consistently decreases. This trend therefore persists across redshifts. It indicates that more massive haloes possess clearer boundaries, and a clearer halo boundary would lead to deeper and narrower splashback features by definition of $\mathcal{D}$ and $\mathcal{W}$.
Based on the power law fitting in Equations 3.1 and 3.1, we add an extra scale factor a=1/(z+1)a=1/{(z+1)} dependence to the width and get the following fitting equations of depth and width of splashback features as a function of both mass and redshift (depth: $\chi_{\nu}^{2}=1.4$; width: $\chi_{\nu}^{2}=2.2$)
@@
Based on the existing fitting equations, we plot the splashback features predicted by halo mass and redshift, which are the dashed lines in the right panels of Figure 3. For haloes with the same mass, the splashback feature is narrower in the past. In addition, the depth does not change significantly with time for a given mass, indicating that the splashback feature was more significant in the past. The narrower and deeper splashback features mean a more rapid change of density of halo in the splashback regions. Part of this trend arises from the growth of haloes: as a halo grows in size with time, by keeping the mass constant, the halo average density is smaller since particles are more spread out. The spreading of particles increases the splashback width. Our analysis in Section 3.3.2 further shows that, at fixed halo mass, halos at earlier times have higher peak heights $\nu$. Since higher $\nu$ halos originate from rarer and more spherically symmetric initial density peaks (Bardeen et al., 1986), it leads to a more coherent gravitational collapse, which sharpens the splashback caustic. The combination of these effects explains why the splashback feature is narrower and deeper at higher redshift. The halo accretion rate also decreases with time (McBride, Fakhouri & Ma, 2009). Slow accretion results in a more gradual transition between virialised and infalling material; thus, the splashback features are smeared out. This trend is further verified quantitatively in the next section through a comparison of the splashback features with different accretion rates.
Qualitatively, we observe that the depth has a smaller dependence on the redshift compared with the width, since the former is approximately constant for different redshifts and the latter is spread out with redshift. This indicates that the splashback width is more affected by the cosmic evolution than the depth is, but overall haloes with the same mass have more distinct splashback features at higher redshifts.
@@ 3.2 Splashback features as a function of halo accretion rate and merger history
Figure 5: Relationships between width ($\mathcal{W}$) and depth ($\mathcal{D}$) of the splashback feature as a function of redshift ($z$) and accretion rate ($\Gamma$). When haloes are stacked by accretion rate, the width increases with time. However, the depth is uncorrelated with redshift and both width and depth appear only weakly dependent on accretion rate. In all panels, points show the data from fits in the simulation, with colour indicating either redshift (left) or accretion rate (right). Haloes are stacked into accretion rate bins between $\Gamma=0$ and $\Gamma=6$ and the median profile is fitted. Top left: Depth of the splashback features as a function of accretion rate for various redshifts. Top right: Depth of the splashback feature as a function of redshift for various accretion rates. There is no clear trend of the depth and width as a function of accretion rate. Bottom left: Width of the splashback features as a function of accretion rate for various redshifts. Bottom right: Width of the splashback feature as a function of redshift for various accretion rates. The black dashed line is fitted with the redshift only, while the black solid line is fitted using the scale factor. There is not much of a trend for the depth and width in terms of halo accretion rate, so we fit a function only in terms of redshift. There is less dependence on redshift when separated by accretion rate compared to the mass bins, as shown in Figure 3, although there is still some dependence, with the depth decreasing significantly between redshifts $0<z<1$ and with the width peaking at redshift $z\sim 1$.
@@
The assembly history of haloes, including the timing and nature of mergers, influences the splashback radius by disrupting the orbits of infalling material. Several studies (e.g. Diemer & Kravtsov, 2014; Adhikari, Dalal & Chamberlain, 2014) show that the splashback feature has a strong dependence on the halo mass accretion rate. Diemer & Kravtsov (2014) notes that high accretion rate haloes with $\Gamma\gtrsim 3$ have a logarithmic slope that reaches approximately 4-4 while lower accretion rate haloes with $\Gamma\lesssim 1$ have logarithmic slopes that reach only 3-3. This observation correlates with the splashback depth defined in our study.
@@
We start with haloes at $z=0$. In Figure 4, the splashback depth gradually increases from 2.4 to 3.0 with the accretion rate from $\Gamma=0$ to $\Gamma=5$. The result matches the observations and conclusions in Diemer & Kravtsov (2014). Although the definition of the splashback depth is not the same as the logarithmic slope minimum in Diemer & Kravtsov (2014), they demonstrate a similar trend. We observe that, at $\Gamma\approx 1$, the splashback depth is approximately $\mathcal{D}=2.4\pm 0.1$; at $\Gamma\approx 3$, the splashback depth is around $\mathcal{D}=2.7\pm 0.1$. This value is consistent with that found in Diemer & Kravtsov (2014). Since Diemer & Kravtsov (2014) did not provide measurements beyond $\Gamma\geq 3$, we extrapolate their reported trend to $\Gamma=5$. Our simulations yield a splashback depth of $3.0\pm 0.6$, consistent with this extrapolation. We attempt to apply the numerical fitting to the splashback depth $\mathcal{D}$ in terms of halo accretion rate. We implement a power law and a linear fitting as below
@@
and they respectively possess reduced $\chi_{\nu}^{2}=0.10$ and $\chi_{\nu}^{2}=0.15$. It is unclear which fitting method better describes the data. In the following section, we show that the splashback depth scales as a power law with peak height. One possible explanation for this scaling is its indirect connection to the halo accretion rate. As shown in Diemer & Kravtsov (2014), the accretion rate increases approximately linearly with peak height, suggesting that the depth may inherit this dependence through the accretion history of the halo.
@@
Therefore, faster accreted haloes have more pronounced halo boundaries and more distinct splashback features. The decrease of the splashback width is also correlated with the results found in Figure 2. Since the halo accretion rate $\Gamma$ increases with the halo mass $M_{200\text{m}}$ and peak height, the power law decrease of the splashback width with accretion rate is consistent with that found in Section 2.2.1 and also the following Section 3.3.2. As haloes accrete matter more rapidly, newly infalling particles pile up more tightly near the first apocentre, leading to a steeper density gradient and a narrower transition zone between the virialised and infalling regions. This results in a narrower splashback width and a more pronounced splashback feature.
In order to check if the splashback feature as a function of accretion rate also persists at higher redshifts, we go back to $z=2$ and plot the depth and width in terms of accretion rate and redshift in Figure 5. At higher redshift, the splashback features tend to be more independent of the halo accretion rate and the previous power law fitting fails. The depth varies between $\mathcal{D}=2.4\sim 2.8$, and the width remains constant between $1\leq\Gamma\leq 6$. At $z\sim 2$ and $\Gamma\gtrsim 4$, the splashback depth decreases with accretion rate. By examining the halo profiles at $z=2$, this can be explained because faster accreting haloes have more turbulent environments (i.e., more density fluctuations at the edges) (Klessen & Hennebelle, 2010), which causes the outskirts past $R_{\rm st}$ to flatten to a lower value at a large radius. Meanwhile, the width at higher redshifts remains insensitive to accretion rate, indicating that the halo boundary remains largely unaffected by accretion rate. Overall, the accretion rate at high redshift affects the halo environment more than the halo internal density structure.
To account for this difference, we examine the halo mass-accretion rate distributions at $z=0$ and $z=2$, which are shown in Figure 6. It shows that, at present, among different accretion bins, the halo mass distributions are different. At $z=0$, the median halo mass in each accretion bin is weakly correlated with accretion rate. In contrast, at $z=2$, the median halo mass remains approximately constant across accretion bins. This may be partially due to the lack of large haloes at $z=2$ making the trend less significant at higher redshift. Therefore, when bootstrapping haloes in each accretion rate bin, there is a stronger dependence on the halo mass at $z=0$ and so higher similarity with splashback features binned by mass in Figure 3. On the other hand, as the mass distributions in each accretion bin are similar at $z=2$, this removes the mass dependence during bootstrapping.
@@
with $\chi_{\nu}^{2}=2.4$. The constant $C$ is the splashback width at $z=0$. More specifically, it is a halo mass accretion rate-dependent variable $C(\Gamma)$ at the present, which is effectively Equations 17 and 19. The width can therefore be fitted as
@@
Unrelaxed haloes that are still adjusting post-merger may exhibit disrupted or less refined splashback features due to chaotic particle orbits compared to relaxed haloes with more stable structures. More et al. (2016) noted that the variation in $\frac{R_{\rm sp}}{R_{\rm 200m}}$ due to the variation in relaxation state is approximately 0.2. Wang, Wang & Mo (2022) also showed that mergers can reduce the splashback radius by up to $10-15\%$ in individual haloes. Therefore, we define the halo merging time in Section 2.2.3 to track this effect. The halo merges when the ratio between the halo’s second most massive subhalo and the most massive subhalo is greater than 0.1. We focus on the haloes at redshift $z=0$. We select haloes with a major merging time less than $z=9$ and and stack the haloes into bins with an interval of $\Delta z=0.5$. In the upper panel of Figure 7, the splashback depth increases before $z=3$, and tends to stability afterwards. This verifies that relaxed haloes which merged at higher redshift tend to have more stable splashback features compared with unrelaxed haloes which merge at $z<3$. The splashback width is relatively constant at $z<6$ due to large error bars. At $z>6$, the splashback width reduces slightly for $\Delta\mathcal{W}\approx 0.1$.
@@
However, the overall density profile is the same for haloes with both recent major mergers and haloes with major mergers that occurred longer ago. This observation is also consistent with the finding of Diemer & Kravtsov (2014), which showed that recent mergers do not alter the stacked halo density profiles when binned by accretion rate. Therefore, the splashback feature depth-to-width ratio is independent of the halo merging time. One thing to note is that these are stacked profiles, but the mergers occur in individual haloes and at only one point on the halo’s edge. Thus, the effects of mergers could be washed out by stacking, which by design removes spherical asymmetries in the halo profiles. It may be possible to align the stacked haloes by merger direction or stack individual objects by solid angle to better capture these effects, but we leave this for future work.
The splashback features demonstrate less dependence on accretion rate and merging time which are short term factors. This shows that instead of being similar to the splashback radius which has a stronger correlation with halo accretion rates, the splashback depth and width depend more on the long term features of haloes. This is possibly due to the depth and width being measures of halo density profile rather than the tracker of particles’ first turnaround at the halo boundary.